Four times the larger number was 11 greater than 7 times the smaller number. If the smaller number was 8 less than the larger number, what was the smaller number?
Answer by mananth(12270)

let larger number be x
smaller number be y
..
4x =7y+11
4x-7y=11...................1
x-y=8......................2
..
multiply equation 2 by -4 and add to equation 1
-4x+4y+4x-7y=-32+11
-3y=-21
y=7
..
x-y =8
x=8+y
x=15